Integrate w.r.t x,
x^3/[(x^2)+1]^3

Grade:12

1 Answers

Vikas TU
10 years ago

x^3/[(x^2)+1]^3.dx

x^2 + 1 = t

diffrntiate both sides

x dx = dt/2

and also, x^2 = t - 1

Now, x^3/[(x^2)+1]^3.dx can be written as {(x)(x^2)/[(x^2)+1]^3.}dx

It will become (t-1)dt/2t3

or (1/2t2)dt - (1/2t3)dt

=Now u can easily integrate and after then replace t by x^2 + 1.

ok

 

 

Approve yarrrrrrrrr

plz!